An instructor gives 4 exams and a final which counts as 2 exams. Find a student's overall average if her exam scores are 62, 83, 97, 90 and an 82 on the final. (round to 2 decimal places.)
Answer #1

Solution:

Given in the question

4 Exams scores are 62,83,97 and 90

Final exam score is 82 which will count as 2

Here we will use weighted average formula

Weighted average = Summation(W*X)/Summation(W) = ((62*1)+(83*1)+(97*1)+(90*1)+(82*2))/(1+1+1+1+2) = (62+83+97+90+164)/6 = 496/6 = 82.67

Student overall average = 82.67